OpenStudy (anonymous):

write the expression using scientific notation (2.5*10^4)(4*10^3) (a) 10*10^7 (b) 10*10^8 (c) 1*10^8 (d) 10*10^12

OpenStudy (anonymous):

(a*x^m)(b*x^n)=(a*b)*10^(m+n)

OpenStudy (anonymous):

yeah umm im not sure that helped.. @minoz

OpenStudy (anonymous):

so did u get ans now?/

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

2.5*4=10 10^4*10^3=10^7 =10*10^7

OpenStudy (anonymous):

oh okay thanks